OpenStudy (anonymous):

-Factor Out The GCF Of The Polynomial x^3-4x^2-21x

OpenStudy (anonymous):

Is there anything besides x that you can factor out?

OpenStudy (anonymous):

What is left is easy to factor. Do the first step by factoring x out.

OpenStudy (anonymous):

I will do the first step for you. \[x^3-4x^2-21x =x( x^2-4x-21)\]

OpenStudy (anonymous):

\[ x^2 - 4 x - 21= (x -?)( x+?) \]

OpenStudy (anonymous):

So How Do You Factor Out The Rest?

OpenStudy (anonymous):

(x-7)(x+3)

OpenStudy (anonymous):

Can You Exsplain How You Got It?

OpenStudy (anonymous):

x1+x2=-(-4) x1.x2=-21

OpenStudy (anonymous):

sum 4 and product 21

OpenStudy (anonymous):

(x -7)(x + 3) = x·x -x·7 + 3·x +-21 First Outer Inner Last = x^2 -4x -21